# Many-Body Localization in the Age of Classical Computing

### v2

@@ VI.3.2 Sensitivity to local perturbations
Figure 9: Parametric motion of eigenvalues $E_{n}(\lambda)$ for the disordered XXZ spin chain. The Hamiltonian reads $H(\lambda)=H_{XXZ}+\lambda\sum_{i=1}^{L}u_{i}\hat{S}^{z}_{i}$, where $H_{XXZ}$ is given by Eq. (7) with $W=0$ and $u_{i}$ are independent random variables distributed uniformly in the interval $[-1,1]$. In ergodic regime, (a), we observe avoided level crossings due to global correlations in the system. In the MBL regime, (b), the regular motion of $E_{n}(\lambda)$ and the presence of avoided crossing indicate the appearance of constants of motion localized in the real space. Reprinted and adapted from [405].
@@
To statistically analyze the properties of the level dynamics, it is convenient to define level velocities $\partial_{\lambda}E_{n}$ [606] and level curvatures, $\partial^{2}_{\lambda}E_{n}$, and to compare their distributions with RMT predictions [725, 453]. The Hellmann–Feynman theorem allows for immediate identification of the level velocities $\partial_{\lambda}E_{n}$ and curvatures $\partial^{2}_{\lambda}E_{n}$ with expressions involving matrix elements of the perturbing operator $\partial_{\lambda}\hat{H}$. The study of level velocities and curvatures reproduces the position of the ETH-MBL crossover identified in exact diagonalization examinations of the disordered spin chains described in the preceding Sections, but, interestingly, certain deviations from the RMT predictions can already be observed deep in the ETH regime [405].
The parametric dependence of the Hamiltonian on $\lambda$ can be introduced with various types of perturbations $\partial_{\lambda}\hat{H}$. Introducing a magnetic flux of strength determined by $\lambda$ is one notable choice. It corresponds to a "twist" of boundary conditions [204, 495]. In that case, the properties of level curvatures can be linked to the notion of Thouless time [655, 14] [^23] The intuition from single-particle physics is that the sensitivity to the ”twist” of boundary conditions determines the extent of delocalization/localization of eigenstates of the system.. Another option is to subject the system to an imaginary magnetic flux [266]. This choice breaks down the hermiticity of $\hat{H}(\lambda)$ and defines a characteristic length scale, which can be identified with localization length in the system. This approach yields results consistent with the exact diagonalization results at the system sizes available to both methods [274, 449].
Investigations of level dynamics under local perturbations such as $\partial_{\lambda}H=S^{z}_{j}$ have been attempted to provide a link into the structure of integrals of motions that arise in the MBL regime. The relatively rare avoided level crossings deep in the MBL regime can be understood as resonances between LIOM configurations [219], which occur on increasing length scales as the disorder strength is decreased towards the ETH-MBL crossover [682, 681].
@@
where $\mathcal{N}$ is the Hilbert space dimension, equals to the norm of an adiabatic gauge potential $\mathcal{A}_{\lambda}$, i.e., the generator, $\mathcal{A}_{\lambda}\ket{n(\lambda)}=i\partial_{\lambda}\ket{n(\lambda)}$, of the adiabatic deformation of the eigenstates. The latter was shown to be a particularly sensitive probe of quantum chaos [463]. Indeed, after a proper regularization of (43) by an introduction of a small energy cutoff, the averaged fidelity $\chi_{\mathrm{av}}$ scales with $L$ exponentially for ergodic systems, polynomially for integrable systems, and it is approximately system-size independent in non-interacting systems (as long as the perturbation $\partial_{\lambda}H$ preserves the class of the model). This hierarchy of scalings motivated [373] to use the fidelity susceptibility as the quantity defining the chaotic behavior, both in quantum system and in their classical limit.
Since the average fidelity susceptibility $\chi_{\mathrm{av}}$ diverges in the XXZ spin chain (7) with $L$ at any finite disorder strength $W$, [572] considered a perturbation $\partial_{\lambda}H=\hat{S}^{z}_{j}$ and computed a typical fidelity susceptibility,
$\chi_{\mathrm{typ}}=\exp\left(\langle\langle\ln(\chi_{n})\rangle\rangle\right),$ (44)
@@
Rescaling the fidelity susceptibility by the Hilbert space dimension $\mathcal{N}$, we obtain a variable $x_{n}=\chi_{n}/\mathcal{N}$, whose exact distribution for the GOE ($\beta=1$) and the GUE ($\beta=2$) is given at $\mathcal{N}\gg 1$ [598] as
@@ VI.3.3 Other approaches and outlook
Studies of ergodicity breaking have also focused on other properties of the matrix elements, among others, on the distribution of the diagonal matrix elements of local observables [138, 349] or on the decay of the difference between diagonal matrix elements in neighboring eigenstates [384]. The latter does not decay with system size $L$ in the MBL regime, in sharp contrast with the exponential in $L$ decay found in the ETH regime [291, 69]. System size drifts towards the ergodic behavior with increasing the system size $L$ are visible in the properties of distributions of matrix elements of local observables as seen, e.g., by [138].
@@
A study [462] of the Binder cumulant $\mathcal{B}$ [76], a quantity proposed as a fine probe for quantum phase transitions, highlights the significance of the finite system size drifts at the ETH-MBL crossover. The Binder cumulant $\mathcal{B}$ is calculated from off-diagonal matrix elements of a local observable ($\hat{S}^{z}_{j}$ operator) between eigenstates of (7) at a fixed finite energy difference $\omega$. Hence, $\mathcal{B}$ probes properties of the system at times which do not scale with $L$, in contrast, e.g., to [387], which focuses at energy scales with the corresponding times of the order of $t_{H}\propto e^{cL}$. The system size and disorder strength behavior of $\mathcal{B}$ does not show any clear signatures of a transition to the MBL phase. However, the observed behavior does not exclude a transition to the MBL phase at sufficiently large $L$ or $W$. Extrapolation of the scaling of $\mathcal{B}$ with $L$ and $W$ allows to obtain a lower bound on the system size $L_{0}$ needed to observe a transition to the MBL phase under an assumption that that the transition occurs at a disorder strength $W_{c}$. For instance, [462] finds that the MBL transition at $W_{c}=10$ would be observable in the Binder cumulant behavior for system sizes larger than $L_{0}\approx 57$.
@@
A different perspective was pursued in studies of a single-particle density matrix, whose matrix elements are defined as $\rho_{ij}=\braket{m}{\hat{c}^{\dagger}_{i}\hat{c}_{j}}{m}$, where |$\ket{m}$ is an eigenstate of a many-body system. Diagonalization of this matrix yields eigenvalues $n_{\alpha}$, interpreted as occupations of single-particle orbitals, which are expected to exhibit a step-like behavior in the MBL regime [57]. Nevertheless, indicators of ergodicity breaking derived in this approach [98, 276, 277] are characterized by finite size drifts similar to the other quantities considered here. Analysis of properties of single-particle orbitals, i.e., the eigenvectors of $\rho_{ij}$, and the related dynamical quantities, such as single-particle Green’s functions, provides a further outlook on the ergodicity and its breakdown in disordered many-body systems [683, 296, 498, 297, 530].
@@
To summarize, matrix elements of local observables provide a deep link between the spectral properties of many-body systems and their dynamics. Hence, understanding the properties of the matrix elements of local observables is vital for assessing ergodicity of the considered system. Focusing on the spectral function $f^{2}(\omega)$, we have shown how features of ergodic and MBL regimes are encoded in the properties of matrix elements of local observables. In particular, the onset of the universal delocalized dynamics at large times manifests itself as a plateau of spectral function at small frequencies, leading to yet another definition of Thouless energy in many-body systems. We have argued that the essential trends observed in the analysis of spectral properties and of eigenstates, which prevent unambiguous conclusions about the fate of the ETH-MBL crossover in the $L\to\infty$ limit, are also present in the properties of matrix elements of local observables.
@@ VII Slow dynamics in disordered many-body systems
This Section concentrates on the time evolution of disordered many-body systems, highlighting the abrupt slowdown of the dynamics in the MBL regime. Focusing on the disordered XXZ model, we discuss the ubiquity of the regime of slow dynamics and outline numerical results for time evolution obtained with sparse matrix time propagation approaches and tensor network algorithms. We show how different scenarios for the fate of the ETH-MBL crossover discussed in Sec. VI translate to the properties of the dynamics in many-body systems. Considering local observables as well as entanglement entropy, we argue that present-day numerical, as well as experimental, time evolution results allow for multiple and mutually inconsistent extrapolations to the limit of large times, $t\to\infty$, and system sizes, $L\to\infty$, relevant for the MBL phase (6).
@@ VII.1.1 Density correlation functions
One approach to probe thermalization in an isolated many-body system is to perform a quench experiment studying the time evolution of a autocorrelation function
$C(t)=\frac{1}{c_{0}}\sum_{j\in\mathcal{J}}\left\langle S^{z}_{j}(t)S^{z}_{j}\right\rangle,$ (46)
@@
where $s_{j}=\pm 1$ is determined by the initial state |$\ket{\vec{\sigma}}$ [^25] In particular, if initial state is taken as the Néel state, $|\ket{\uparrow\downarrow\uparrow\downarrow\ldots}$, the autocorrelation function $C(t)$ is sometimes called an imbalance. We do not follow this distinction here as the time evolution of the Néel state follows the same patterns as for typical product states with energies close to the middle of the many-body spectrum.. The ETH is valid (4) in the ergodic regime. For the disordered XXZ spin chain (and the $J_{1}$-$J_{2}$ model), the microcanonical average of $S^{z}_{j}$ vanishes in the large system size limit, c.f., Fig. 8(c). Therefore, the autocorrelation function $C(t)$ decays to zero in the ergodic regime, and the information about the initial spin configuration |$\ket{\vec{\sigma}}$ is scrambled during the system’s time evolution. In contrast, in the MBL regime, the approximate LIOMs preserve some fraction of the information about the initial spin configuration |$\ket{\vec{\sigma}}$ indefinitely, leading to a non-vanishing autocorrelation function $C(t)$ in the $t\to\infty$ limit. The memory of the initial state retained at sufficiently strong disorder in many-body systems has been the basis of experimental observations of the MBL regime, see e.g. [561].
The decay of the autocorrelation function $C(t)$ to its long-time average contains information about the dynamics in the system. Diffusive dynamics is one possible scenario in systems with a conserved quantity, e.g., conservation of the total magnetization $\sum_{j}S^{z}_{j}$ in the considered disordered spin chains may lead to diffusive spin dynamics. Then, at sufficiently large time and length scales, an effective hydrodynamic description of the dynamics is possible [304], which predicts a decay $C(t)\propto t^{-1/2}$ for 1D systems. Such a hydrodynamic behavior is inevitably accompanied by scaling corrections significant at intermediate time and length scales. The scaling corrections may be constrained due to additional symmetries in the system [143] and have to be taken into account when assessing the class of dynamics of the considered system [415]. Another scenario for systems with a conserved density is a subdiffusion with dynamical exponent $z>2$, which leads to a hydrodynamic description with an algebraic decay of the density autocorrelation function $C(t)\propto t^{-1/z}$ [165].
Figure 10: Time evolution of density autocorrelation function $C(t)$ in disordered XXZ spin chain (7). (a) For $W=4$, the interactions ($\Delta=1$) induce a slow decay of $C(t)$ at long times, which continues up to approximately Heisenberg time $t_{H}$ (denoted by the red squares). For $\Delta=0$, in the Anderson localized phase, $C(t)$ does not decay after the initial oscillations. (b) For disorder strength $W=3$, a strong decay of $C(t)$ is observed as the XXZ spin chain is in the crossover regime. (c) For $W=5$ the system is in the MBL regime $W>W^{*}(L)$ for $12\leq L\leq 20$. The decay of $C(t)$ is persistent, but very slow, and an understanding of whether the system thermalizes in the $L,t\to\infty$ limit is missing. Reprinted and adapted from [604].
@@
In the presence of interactions, $\Delta=1$, and at the same disorder strength $W=4$, the autocorrelation function $C(t)$ slowly decays in time and ceases to decay only around the Heisenberg time $t_{H}\propto e^{cL}$, see Fig. 10(a). The persistent decline of $C(t)$ is the effect of interactions in the system, and, despite being slow, the decay may have a strong impact on the fate of the system in the asymptotic $L,t\to\infty$ limit. The system’s behavior illustrates this for $W=3$, shown in Fig. 10(b). At considered system sizes $12\leq L\leq 20$, the value of $C(t)$ at any $t$ is no smaller than $0.4$, far from the ETH prediction $C(t)\stackrel{{\scriptstyle t\to\infty}}{{\longrightarrow}}0$ (relevant for $L\gg 1$). Hence, we observe a clear breakdown of ergodicity at considered $L$, which is consistent with the value of the average gap ratio at $W=3$, c.f. Sec. VI.1, according to which the system is in the crossover regime $W^{T}(L)<W<W^{*}(L)$. Nevertheless, the saturation values of $C(t)$ are, to a good approximation, decreasing proportionally to $L$. Extrapolation of this trend indicates that in the limit of large $L$ and $t$, the autocorrelation function $C(t)$ decays to 00, and the system thermalizes.
@@
The observation that power-law fits accurately describe results in the crossover regime of disordered XXZ spin chain, at disorder strengths roughly fulfilling $W^{T}(L)<W<W^{*}(L)$, was attributed, by the early works [239, 241, 7, 461], to heterogeneities in disordered systems. Spatially uncorrelated disorder necessarily leads to Griffiths regions, where the disorder is locally stronger. Such regions play an essential role in the physics of quantum phase transitions [684], and may act as bottlenecks hindering the thermalization of the surrounding spins and generating the slow power-law decays [394, 388] of local correlators observed experimentally by [399]. The influence of Griffiths regions on the stability of the MBL phase will be revisited in Sec. IX. In passing, we note that the power-law decays were also observed in the dynamics of survival probabilities [642], which are discussed in more details in Sec. VII.3.5.
@@
The extrapolation of the results to the $L,t\to\infty$ limit is less straightforward when disorder strength $W$ in an interacting system is increased beyond $W\gtrapprox 3$. The power-law decay of $C(t)$, which accurately describes data at $W=3$, is no longer a good fit for $W=5$ and system sizes up to $L=20$. Indeed, as visible in Fig. 10(c), the decay is slower than logarithmic in time. Subsequent studies realized that an increase of the system size beyond the capabilities of ED, possible due to Chebyshev time evolution method [55] (up to $L=24$) or TEBD algorithm [174] (up to $L=100$), increases the prominence of the power-law decays in XXZ spin chains at disorder strengths bigger than $W^{*}(L=20)\approx 4$. The tensor network studies [174, 111], employing various types of threshold on the exponent $\beta$ governing the observed power-law decay of $C(t)$ (see also [492, 175]), interpreted the observed behavior as a manifestation of the MBL transition occurring at disorder strengths above the early ED estimates [393], c.f. [443]. Similar approaches in which the average in (47) is taken over specific initial states [497] show that the onset of the MBL regime is dependent on the energy of the initial product state, which was interpreted as a premise for many-body mobility edge in disordered spin chains [439, 110].
@@
Figure 11: Persistent slow dynamics in disordered XXZ spin chain (7). (a) Deep in the MBL regime, at $W=8$, the density autocorrelation function $C(t)$, calculated with TDVP algorith, slowly decays, the decay can be fitted with a power-law $C(t)\propto t^{-\beta}$, with $\beta\approx 10^{-3}$. (b) The dynamical exponent $\beta$ from: [388] [Luitz ’17], governing the power-law decay of $\left\langle S^{z}_{j}(t)S^{z}_{j}\right\rangle$ obtained with typicality approach for $L=24$; [174] [Doggen ’18] governing the power-law decay of $C(t)$ for initial Néel state, calculated for $L=100$ with TDVP; [111] for decay of $C(t)$ for initial Néel state, calculated with Chebyshev time propagation for $L=26$ (see Sec. V.2) [Chanda ’20; Cheby], and with TDVP for $L=50$ [Chanda ’20; TDVP]; [604] [Sierant ’22] for decay of $C(t)$ for initial Néel state calculated with TDVP for $L=50,200$. We observe an exponential decrease $\beta\propto e^{-\gamma W}$, where $\gamma=0.8$ is a constant. Panel (a) is reprinted and adapted from [604].

### v3

@@ VI.3.2 Sensitivity to local perturbations
Figure 9: Parametric motion of eigenvalues $E_{n}(\lambda)$ for the disordered XXZ spin chain. The Hamiltonian reads $\hat{H}(\lambda)=H_{XXZ}+\lambda\sum_{i=1}^{L}u_{i}\hat{S}^{z}_{i}$, where $H_{XXZ}$ is given by Eq. (7) with $W=0$ and $u_{i}$ are independent random variables distributed uniformly in the interval $[-1,1]$. In ergodic regime, (a), we observe avoided level crossings due to global correlations in the system. In the MBL regime, (b), the regular motion of $E_{n}(\lambda)$ and the presence of avoided crossing indicate the appearance of constants of motion localized in the real space. Reprinted and adapted from [441].
@@
To statistically analyze the properties of the level dynamics, it is convenient to define level velocities, $\partial_{\lambda}E_{n}$ [655], and level curvatures, $\partial^{2}_{\lambda}E_{n}$, and to compare their distributions with RMT predictions [781, 491]. The Hellmann–Feynman theorem allows for immediate identification of the level velocities $\partial_{\lambda}E_{n}$ and curvatures $\partial^{2}_{\lambda}E_{n}$ with expressions involving matrix elements of the perturbing operator $\partial_{\lambda}\hat{H}$. The study of level velocities and curvatures reproduces the position of the ETH-MBL crossover identified in exact diagonalization examinations of the disordered spin chains described in the preceding Sections, but, interestingly, certain deviations from the RMT predictions can already be observed deep in the ETH regime [441].
The parametric dependence of the Hamiltonian on $\lambda$ can be introduced with various types of perturbations $\partial_{\lambda}\hat{H}$. Introducing a magnetic flux of strength determined by $\lambda$ is one notable choice. It corresponds to a "twist" of boundary conditions [216, 535]. In that case, the properties of level curvatures can be linked to the notion of Thouless time [707, 15] [^33] The intuition from single-particle physics is that the sensitivity to the ”twist” of boundary conditions determines the extent of delocalization/localization of eigenstates of the system.. Another option is to subject the system to an imaginary magnetic flux [285], a method that has been explored for interacting models in the context of MBL [280]. This choice breaks down the hermiticity of $\hat{H}(\lambda)$ and defines a characteristic length scale, which can be identified with localization length in the system. This approach yields results consistent with the exact diagonalization results at the system sizes available to both methods [293, 487].
Investigations of level dynamics under local perturbations, such as $\partial_{\lambda}H=\hat{S}^{z}_{j}$, have been attempted to provide a link into the structure of integrals of motions that arise in the MBL regime. The relatively rare avoided level crossings deep in the MBL regime can be understood as resonances between LIOM configurations [234], which occur on increasing length scales as the disorder strength is decreased towards the ETH-MBL crossover [737, 736].
@@
where $\mathcal{N}$ is the Hilbert space dimension, equals to the norm of an adiabatic gauge potential $\mathcal{A}_{\lambda}$, i.e., the generator, $\mathcal{A}_{\lambda}\ket{n(\lambda)}=i\partial_{\lambda}\ket{n(\lambda)}$, of the adiabatic deformation of the eigenstates. The latter was shown to be a particularly sensitive probe of quantum chaos [501]. Denoting $\omega_{mn}=E_{m}-E_{n}$ and regularizing the energy denominator in (42) as $1/\omega_{mn}^{2}\to\omega_{mn}^{2}/(\omega_{mn}^{2}+\mu^{2})^{2}$, where $\mu$ is a small energy cut-off, (43) can be rewritten as
[…]
The latter formula relies on the form of the spectral function $f^{2}(\omega)$ (41) of the operator $\hat{A}=\partial_{\lambda}H$ with $\mathcal{M}_{E_{c}}$ containing the full spectrum of the system and with the contribution $C_{0}\delta(\omega)$ stemming from $m=n$ in (41) subtracted. When the energy cut-off is taken to be of the order of inverse level spacing, $\mu\propto\omega_{\mathrm{H}}\propto 2^{-L}$, the average fidelity susceptibility allows for distinguishing different dynamical regimes. For ergodic system, the spectral function $f^{2}(\omega)$ has a plateau at $\omega_{\mathrm{H}}<\omega<\omega_{\mathrm{Th}}$, which, through (44), implies an exponential scaling $\chi_{\mathrm{av}}\propto 2^{L}$. Vanishing of the spectral function at small frequencies for integrable models results in an algebraic scaling of $\chi_{\mathrm{av}}$ with $L$, while for non-interacting models, $\chi_{\mathrm{av}}$ is approximately system-size independent (as long as the perturbation $\partial_{\lambda}H$ preserves the class of the model). This hierarchy of scalings motivated [408] to propose the fidelity susceptibility as the quantity defining the chaotic behavior, both in quantum systems and in their classical limit. This classification distinguishes also a maximally chaotic regime, in which the sensitivity of eigenstates to the perturbation $\partial_{\lambda}H$ is maximal, $\chi_{\mathrm{av}}\propto 4^{L}$. In disordered spin chains, the maximally chaotic regime is found at disorder strengths $W$ above the ergodic regime when the slow down of the dynamics results in the behavior $f^{2}(\omega)\propto\omega^{-\alpha}$ with $\alpha\to 1$.
@@
Rescaling the fidelity susceptibility (42) by the Hilbert space dimension $\mathcal{N}$, we obtain a variable $x_{n}=\chi_{n}/\mathcal{N}$, whose exact distribution for the GOE ($\beta=1$) and the GUE ($\beta=2$) is given at $\mathcal{N}\gg 1$ [647] as
@@ VI.3.3 Other approaches and outlook
Studies of ergodicity breaking have also focused on other properties of the matrix elements, among others, on the distribution of the diagonal matrix elements of local observables [143, 378] or on the decay of the difference between diagonal matrix elements in neighboring eigenstates [420]. The latter does not decay with system size $L$ in the MBL regime, in sharp contrast to the exponential in $L$ decay found in the ETH regime [72], which also differs from the algebraic decay of diagonal matrix element fluctuations in integrable systems [315, 785]. System size drifts towards the ergodic behavior with increasing the system size $L$ are visible in the properties of distributions of matrix elements of local observables as seen, e.g., by [143].
@@
A study [500] of the Binder cumulant $\mathcal{B}$ [79], a quantity proposed as a fine probe for quantum phase transitions, highlights the significance of the finite system size drifts at the ETH-MBL crossover. The Binder cumulant $\mathcal{B}$ was calculated from the off-diagonal matrix elements of a local observable ($\hat{S}^{z}_{j}$ operator) between eigenstates of (7) at a fixed finite energy difference $\omega$. Hence, $\mathcal{B}$ probes properties of the system at times which do not scale with $L$, in contrast, e.g., to [423], which focuses at energy scales with the corresponding times of the order of $t_{H}\propto e^{cL}$. The system size and disorder strength behavior of $\mathcal{B}$ does not show any clear signatures of a transition to the MBL phase. However, the observed behavior does not exclude a transition to the MBL phase at sufficiently large $L$ or $W$. Extrapolation of the scaling of $\mathcal{B}$ with $L$ and $W$ allows to obtain a lower bound on the system size $L_{0}$ needed to observe a transition to the MBL phase under an assumption that the transition occurs at a disorder strength $W_{c}$. For instance, [500] finds that the MBL transition at $W_{c}=10$ would be observable in the Binder cumulant behavior for system sizes larger than $L_{0}\approx 57$.
@@
A different perspective was pursued in studies of a single-particle density matrix, whose matrix elements are defined as $\rho_{ij}=\braket{m}{\hat{c}^{\dagger}_{i}\hat{c}_{j}}{m}$, where |$\ket{m}$ is an eigenstate of a many-body system. Diagonalization of this matrix yields eigenvalues $n_{\alpha}$, interpreted as occupations of single-particle orbitals, which are expected to exhibit a step-like behavior in the MBL regime [60]. Nevertheless, indicators of ergodicity breaking derived in this approach [101, 295, 296] are characterized by finite size drifts similar to the other quantities considered here. Analysis of properties of single-particle orbitals, i.e., the eigenvectors of $\rho_{ij}$, and the related dynamical quantities, such as the single-particle Green’s functions, provides a further outlook on the ergodicity and its breakdown in disordered many-body systems [738, 321, 539, 322, 576].
@@
To summarize, matrix elements of local observables provide a deep link between the spectral properties of many-body systems and their dynamics. Hence, understanding the properties of the matrix elements of local observables is vital for assessing the ergodicity of the considered system. Focusing on the spectral function $f^{2}(\omega)$, we have shown how features of ergodic and MBL regimes are encoded in the properties of matrix elements of local observables. In particular, the onset of the universal delocalization dynamics at large times manifests itself as a plateau of spectral function at small frequencies, leading to yet another definition of Thouless energy in many-body systems. We have argued that the essential trends observed in the analysis of spectral properties and of eigenstates, which prevent unambiguous conclusions about the fate of the ETH-MBL crossover in the $L\to\infty$ limit, are also present in the properties of matrix elements of local observables.
@@ VII Slow dynamics in disordered many-body systems
This Section concentrates on the time evolution of disordered many-body systems, highlighting the abrupt slowdown of the dynamics in the MBL regime. Focusing on the disordered XXZ model, we discuss the ubiquity of the regime of slow dynamics and outline numerical results for time evolution obtained with sparse matrix time propagation approaches and tensor network algorithms. We show how different scenarios for the fate of the ETH-MBL crossover discussed in Sec. VI translate to the properties of the dynamics in many-body systems. Considering local observables as well as entanglement entropy, we argue that the present-day numerical, as well as experimental, time evolution results allow for multiple and mutually inconsistent extrapolations to the limit of large times, $t\to\infty$, and system sizes, $L\to\infty$, relevant for the MBL phase (6).
@@ VII.1.1 Density correlation functions
One approach to probe thermalization in an isolated many-body system is to perform a quench experiment studying the time evolution of an autocorrelation function
$C(t)=\frac{1}{c_{0}}\sum_{j\in\mathcal{J}}\left\langle\hat{S}^{z}_{j}(t)\hat{S}^{z}_{j}\right\rangle,$ (46)
@@
[…]
The decay of the autocorrelation function $C(t)$ to its long-time average contains information about the dynamics in the system. Diffusive dynamics is one possible scenario in systems with a conserved quantity, e.g., conservation of the total magnetization $\sum_{j}\hat{S}^{z}_{j}$ in the considered disordered spin chains may lead to diffusive spin dynamics. Then, at sufficiently large time and length scales, an effective hydrodynamic description of the dynamics is possible [330], which predicts a decay $C(t)\propto t^{-1/2}$ for 1D systems. Such a hydrodynamic behavior is inevitably accompanied by scaling corrections significant at intermediate time and length scales. The scaling corrections may be constrained due to additional symmetries in the system [148] and have to be taken into account when assessing the class of dynamics of the considered system [451]. Another scenario for systems with a conserved density is a subdiffusion with dynamical exponent $z>2$, which leads to a hydrodynamic description with an algebraic decay of the density autocorrelation function, $C(t)\propto t^{-1/z}$ [173].
Figure 10: Time evolution of density autocorrelation function $C(t)$ in the disordered XXZ spin-1/2 chain (7). (a) For $W=4$, the interactions ($\Delta=1$) induce a slow decay of $C(t)$ at long times, which continues approximately up to the Heisenberg time $t_{H}$ (denoted by the red squares). For $\Delta=0$, in the Anderson localized phase, $C(t)$ does not decay after the initial oscillations. (b) For disorder strength $W=3$, a strong decay of $C(t)$ is observed as the disordered XXZ spin-1/2 chain is in the crossover regime. (c) For $W=5$ the system is in the MBL regime, $W>W^{*}(L)$ for $12\leq L\leq 20$. The decay of $C(t)$ is persistent, but very slow, and an understanding of whether the system thermalizes in the $L,t\to\infty$ limit is missing. Reprinted and adapted from [653].
@@
In the presence of interactions, $\Delta=1$, and at the same disorder strength $W=4$, the autocorrelation function $C(t)$ slowly decays in time and ceases to decay only at long times approximately given by the Heisenberg time $t_{H}\propto e^{cL}$, see Fig. 10(a). The persistent decline of $C(t)$ is the effect of interactions in the system, and, despite being slow, the decay may have a strong impact on the fate of the system in the asymptotic $L,t\to\infty$ limit. The system’s behavior illustrates this for $W=3$, shown in Fig. 10(b). At considered system sizes $12\leq L\leq 20$, the value of $C(t)$ at any $t$ is no smaller than $0.4$, far from the ETH prediction $C(t)\stackrel{{\scriptstyle t\to\infty}}{{\longrightarrow}}0$ (relevant for $L\gg 1$). Hence, we observe a clear breakdown of ergodicity at considered $L$, which is consistent with the value of the average gap ratio at $W=3$, c.f. Sec. VI.1, according to which the system is in the crossover regime $W^{T}(L)<W<W^{*}(L)$. Nevertheless, the saturation values of $C(t)$ are, to a good approximation, decreasing proportionally to $L$. Extrapolation of this trend indicates that in the limit of large $L$ and $t$, the autocorrelation function $C(t)$ decays to 00, and the system thermalizes.
@@
The observation that power-law fits accurately describe the results in the crossover regime of the disordered XXZ spin chain, at disorder strengths roughly fulfilling $W^{T}(L)<W<W^{*}(L)$, was attributed, by the early works [255, 257, 7, 499], to heterogeneities in disordered systems. Spatially uncorrelated disorder necessarily leads to Griffiths regions, where the disorder is locally stronger. Such regions play an essential role in the physics of quantum phase transitions [739], and may act as bottlenecks hindering the thermalization of the surrounding spins and generating the slow power-law decays [431, 424] of local correlators observed experimentally by [435]. The influence of Griffiths regions on the stability of the MBL phase will be revisited in Sec. IX. In passing, we note that the power-law decays were also observed in the dynamics of survival probabilities [694], which are discussed in more detail in Sec. VII.3.5. Resonant regions, intimately connected to disorder fluctuations, play also an important role in the mathematical analysis of strongly disordered quantum Ising chains [573]. The demonstration of the MBL in that work is limited only to rare realizations of the 1D chain, indicating the absence of diffusion at strong disorder, but allowing for subdiffusive behavior.
@@
The extrapolation of the results to the $L,t\to\infty$ limit is less straightforward when the disorder strength $W$ in an interacting system is increased beyond $W\gtrapprox 3$. The power-law decay of $C(t)$, which accurately describes the results at $W=3$, is no longer a good fit for $W=5$ and system sizes up to $L=20$. Indeed, as visible in Fig. 10(c), the decay is slower than logarithmic in time. Subsequent studies realized that an increase of the system size beyond the capabilities of ED, possible due to Chebyshev time evolution method [58] (up to $L=24$) or TEBD algorithm [183] (up to $L=100$), increases the prominence of the power-law decays in the XXZ spin chains at disorder strengths larger than $W^{*}(L=20)\approx 4$. The tensor network studies [183, 115], employing various types of threshold on the exponent $\beta$ governing the observed power-law decay of $C(t)$ (see also [532, 184]), interpreted the observed behavior as a manifestation of the MBL transition occurring at disorder strengths above the early ED estimates [430], c.f. [481]. Similar approaches in which the average in (47) is taken over specific initial states [538] show that the onset of the MBL regime is dependent on the energy of the initial product state, which was interpreted as a premise for many-body mobility edge in disordered spin chains [477, 114].
@@
Figure 11: Persistent slow dynamics in the disordered XXZ spin-1/2 chain (7). (a) Deep in the MBL regime, at $W=8$, the density autocorrelation function $C(t)$, calculated with TDVP algorithm, slowly decays, the decay can be fitted with a power-law $C(t)\propto t^{-\beta}$, with $\beta\approx 10^{-3}$. (b) The dynamical exponent $\beta$ from: [424] [Luitz ’17], governing the power-law decay of $\left\langle\hat{S}^{z}_{j}(t)\hat{S}^{z}_{j}\right\rangle$ obtained with typicality approach for $L=24$; [183] [Doggen ’18] governing the power-law decay of $C(t)$ for initial Néel state, calculated for $L=100$ with TDVP; [115] for the decay of $C(t)$ for initial Néel state, calculated with Chebyshev time propagation for $L=26$ (see Sec. V.2) [Chanda ’20; Cheby], and with TDVP for $L=50$ [Chanda ’20; TDVP]; [653] [Sierant ’22] for the decay of $C(t)$ for initial Néel state calculated with TDVP for $L=50,200$. We observe an exponential decrease $\beta\propto e^{-\gamma W}$, where $\gamma=0.8$ is a constant. Panel (a) is reprinted and adapted from [653].